OpenStudy (anonymous):

convert to scientific notation 21,000,000,000

OpenStudy (anonymous):

2.1 * 10^{x} Now just replace x with however many places you moved the decimal place.

OpenStudy (anonymous):

Do you see what I mean? Scientific notation is very useful and used in all the sciences, so it's important to understand how it works. \[\Large 0.00031 = 3.1 \times {10^{ - 4}}\]The decimal has moved 4 places to the right (note the negative).\[\Large 15,000 = 1.5 \times {10^4}\]The decimal has moved 4 places to the left. Note that on a whole number, even if the decimal is not shown, it is considered to be following the last digit since 15,000 = 15,000.000 (etc.) Understand? =)